OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 A -7 B 7 C 9 D -9

OpenStudy (anonymous):

I don't know how to do it that's why I need the answers

OpenStudy (anonymous):

Please read this: It will be better for you: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Lame

OpenStudy (anonymous):

I won't give you answer but I can give you hints on how to solve this.

OpenStudy (anonymous):

5x + 4y = 1..................................1 4x + 3y = -1.......................2 From equation 1 5x=1-4y x=1-4y/5 now put into 2nd equ then get value of y . After getting value of y then put into 1st equ to get value of x.

OpenStudy (anonymous):

Example: 5x+y=31 [1] x-2y=-7 [2] rearrange [1] to make y the subject. y = 31 - 5x rearrange [2] to make y the subject. -2y = -7 - x 2y = 7 + x y = (7 + x)/2 Now equate your two expressions for y. 31 - 5x = (7+x)/2 2(31 - 5x) = 7 + x 62 - 10x = 7 + x 55 = 11x therefore x = 5 Now sub in x = 5 into [1] or [2] (doesn't matter) [2] -> (5)-2y = -7 -2y = -12 2y = 12 y = 6 So there you have it... x = 5, y = 6. Now follow my approach on your problem

OpenStudy (anonymous):

Another approach could be the following: If you multiply the 1st equation by 4 and the second equation by 5, you'll get two equations that both have 20x in them. You can isolate 20X on both and then combine to get a single equation with only Y as the unknown (which is what you're solving for)

OpenStudy (anonymous):

do you need help